# bibliography.bib

### v1

@@ Appendix G Proofs for Weight Sharing
To pre-process the MNIST dataset, we first create a single $D_{32}$ steerable convolution layer with kernel size equal to the images’ size, one input channel and $6400100\times|D_{32}|=$ 6400 output channels. In particular, under the steerable CNNs framework, we use $100$ copies of the regular representation of the group $D_{32}$ as output feature type. Alternatively, in the group-convolution framework, this steerable convolution layer corresponds to lifting convolution with $100$ output channels, mapping a $1$-channel scalar image to a $100$-channels signal over the whole $\mathbb{R}^{2}\rtimes D_{32}$ group; because the kernel size is as large as the input image (because we use no padding), the spatial resolution of the output of the convolution is a single pixel, leaving only a feature vector over $G=D_{32}$.
@@
Our equivariant networks consist of a linear layer (i.e. a group convolution), followed by a ReLU activation and a final pooling and invariant linear layer as in Eq. 3. All models are trained using the Adam optimizer with a learning rate of $21e-$ 2 for $30$ epochs. We use the MNIST12K dataset, which has 12K images in the training split and 50K in the test split.

### v2

@@ Appendix G Proofs for Weight Sharing
Figure 5: Graphical representation of the equivariant linear projection used to preprocess the image data. An example image is projected with a single filter rotated $8$ times and mirrored and rotated $8$ more times. The output is a $8\cdot 2=16$ dimensional vector representing a signal over $G=D_{8}$. A rotation or mirroring of the input image results in a periodic shift or permutation of the output channels.
To pre-process the MNIST and the CIFAR10 datasets, we first create a single $D_{32}$ steerable convolution layer with kernel size equal to the images’ size, one input channel and $6400100\times|D_{32}|=$ 6400 output channels. In particular, under the steerable CNNs framework, we use $100$ copies of the regular representation of the group $D_{32}$ as output feature type. Alternatively, in the group-convolution framework, this steerable convolution layer corresponds to lifting convolution with $100$ output channels, mapping a $1$-channel scalar image to a $100$-channels signal over the whole $\mathbb{R}^{2}\rtimes D_{32}$ group; because the kernel size is as large as the input image (because we use no padding), the spatial resolution of the output of the convolution is a single pixel, leaving only a feature vector over $G=D_{32}$.
@@
Fig. 5 shows an example of an input image projected with a single filter rotated $8$ times and mirrored and rotated $8$ more times (i.e. encoded via a $G=D_{8}$ steerable convolution). The output is a $8\cdot 2=16$ dimensional vector representing a signal over $G=D_{8}$. A rotation or mirroring of the input image results in a periodic shift or permutation of the output channels.
@@
Our equivariant networks consist of a linear layer (i.e. a group convolution), followed by a ReLU activation and a final pooling and invariant linear layer as in eq. 3. All MNIST models are trained using the Adam optimizer with a learning rate of $21e-$ 2 for $30$ epochs, while the CIFAR10 models are trained using with a learning rate of $31e-$ 3 for $50$ epochs. Precisely, we use the MNIST12K dataset, which has 12K images in the training split and 50K in the test split.
@@
Figure 6: Subset of Fig. 2(c) focusing on the models training with the largest dataset size $m=25600$. The models using the smallest equivariance groups show increased norms.